OpenStudy (anonymous):

How many possible solutions are there for these equations? x - 3y = -3, -2x + 6y = 12

OpenStudy (mathmale):

Look carefully at the two equations. Is one of them a multiple of the other? If so, demonstrate how the 2nd is related to the first.

OpenStudy (solomonzelman):

x - 3y = -3 -2x + 6y = 12 divide this equation (each thing in it) by -2 what do you get ?